# Prosperity before Collapse: How Far Can Off-Policy RL Reach with Stale Data on LLMs?

### v1

@@ 4 Prosperity before Collapse: Stale data contain enough training information in RL on LLMs
Pivotal token masking by $\epsilon$-clipping when training with stale data. As also discussed in recent work (Su et al., 2025; Chen et al., 2025), $\epsilon$-clipping may inadvertently mask important tokens, preventing them from contributing useful training signals. We extend this observation to the asynchronous setting and show that the problem becomes substantially more severe when training with stale data, since larger staleness induces a greater mismatch between the behavior and target policies. As illustrated in Figure 4(a), the clipping ratio increases sharply under large staleness ($s=256$), while remaining negligible in the on-policy baseline.
@@
To better understand this phenomenon, we conduct a quantitative analysis on 90 million training tokens collected during Qwen2.5-Math-7B training with staleness $256$. Specifically, we gather all training tokens generated between 800 and 1200 model updates, ensuring the model is already in a stable training phase but before convergence. Figure 4(b) shows a clear trend: as |r1||r-1| increases, the average token entropy also rises. This indicates that $\epsilon$-clipping disproportionately prunes high-entropy tokens, which are typically the most informative for model improvement (Wang et al., 2025a; Gao et al., 2025; Cui et al., 2025). Consequently, clipping under stale data leads to degraded performance.
@@ 5 $\boldsymbol{M}_{j}\leftarrow\texttt{False}$;  $\mathcal{T}\leftarrow\mathcal{T}\setminus\{j\}$;
$\mathcal{J}_{\text{M2PO}}(\theta)=\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\boldsymbol{M}_{i,t}\,\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}\,A_{i,t},\qquad\boldsymbol{M}_{i,t}\in\{0,1\},$ (5)
@@ Acknowledgment
We would like to thank Cheng Luo, Xinyu Yang, Ranajoy Sadhukhan, Xuesheng Liu, Yongji Wu for providing us constructive feedback on our paper and the computing resources of NVIDIA. This work is supported in part by the grants NSF CCF-2504353 to B. Chen. This work is also partially supported by Google Research Award, Amazon Research Award, Intel, Li Auto, Moffett AI, and CMU CyLab Seed funding. We are also grateful to BitDeer AI Research for providing GPU resources Any opinions, findings, and conclusions or recommendations expressed are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ 11 Additional Experiments
Commonly clipped tokens in GRPO. Figure 10 shows the specific tokens that are most frequently clipped by $\epsilon$-clipping. The word cloud of commonly clipped token is highly aligned with high-entropy tokens shown in (Wang et al., 2025a): these tokens are not random or unimportant, but rather belong to the most semantically and structurally critical elements in reasoning traces. Many of them (e.g., First, simplify, determine, To def, Thus, verify, break) are precisely the high-entropy “pivotal tokens” that initiate, connect, or conclude key reasoning steps. Others (e.g., assistant, user, code markers like ### or \$\$) serve as structural anchors in the dialogue or mathematical formatting. This observation aligns with Figure 4(b): as the importance weight ratio |r1||r-1| grows, clipped tokens tend to exhibit higher entropy.

### v2

@@ 4 Prosperity before Collapse: Stale Data Contain Enough Training Information in RL on LLMs
Pivotal token masking by $\epsilon$-clipping when training with stale data. As also discussed in recent work (Su et al., 2025; Chen et al., 2025), $\epsilon$-clipping may inadvertently mask important tokens, preventing them from contributing useful training signals. We extend this observation to the asynchronous setting and show that the problem becomes substantially more severe when training with stale data, since larger staleness induces a greater mismatch between the behavior and target policies. As illustrated in Figure 4(a), the clipping ratio increases sharply under large staleness ($s=256$), while remaining much lower in the on-policy baseline.
@@
To better understand this phenomenon, we conduct a quantitative analysis on 90 million training tokens collected during Qwen2.5-Math-7B training with staleness $256$. Specifically, we gather all training tokens generated between 800 and 1200 model updates, ensuring the model is already in a stable training phase but before convergence. Figure 4(b) shows a clear trend: as |r1||r-1| increases, the average token entropy also rises. This indicates that $\epsilon$-clipping disproportionately prunes high-entropy tokens, which are typically the most informative for model improvement. Consequently, clipping under stale data leads to degraded performance.
@@ 5 $\boldsymbol{M}_{j}\leftarrow\texttt{False}$;  $\mathcal{T}\leftarrow\mathcal{T}\setminus\{j\}$;
$\mathcal{J}_{\text{M2PO}}(\theta)=\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\boldsymbol{M}_{i,t}\,\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{behav}}(o_{i}|q)}\,A_{i,t},\qquad\boldsymbol{M}_{i,t}\in\{0,1\},$ (5)
@@ Acknowledgment
We would like to thank Cheng Luo, Xinyu Yang, Ranajoy Sadhukhan, Xuesheng Liu, Yongji Wu for providing us constructive feedback on our paper and the computing resources of NVIDIA. This work is supported in part by the grants NSF CCF-2504353 to B. Chen. This work is also partially supported by Google Research Award, Amazon Research Award, Intel, Li Auto, Moffett AI, and CMU CyLab Seed funding. We are also grateful to BitDeer AI Research for providing GPU resources. Any opinions, findings, and conclusions or recommendations expressed are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ 11 Additional Experiments
Commonly clipped tokens in GRPO. Figure 10 shows the specific tokens that are most frequently clipped by $\epsilon$-clipping. The word cloud of commonly clipped tokens is highly aligned with high-entropy tokens shown in (Wang et al., 2025a): these tokens are not random or unimportant, but rather belong to the most semantically and structurally critical elements in reasoning traces. Many of them (e.g., First, simplify, determine, To def, Thus, verify, break) are precisely the high-entropy “pivotal tokens” that initiate, connect, or conclude key reasoning steps. Others (e.g., assistant, user, code markers like ### or \$\$) serve as structural anchors in the dialogue or mathematical formatting. This observation aligns with Figure 4(b): as the importance weight ratio |r1||r-1| grows, clipped tokens tend to exhibit higher entropy.